# Automated Testing Method for Text-to-Image Software

### v2

@@ 5.5. Threats to Validity
Test subject. The selection of T2I software is one of the vital threats to validity. With the rapid development of T2I software, a diverse array of such software is continually emerging, each with varying engines and mechanisms. To alleviate this threat, we employ four commonly used and popular T2I software powered by different core engines.
@@
Dataset Selection The selection of the source dataset is another threat that comes from. Due to the fact that not all training and validation datasets used by various software are publicly available, there is a potential issue of data domain unfairness. To alleviate this threat, we select the most commonly used and classic open-source dataset MS-COCO as our seed collection to ensure a level of fairness in the experiments.
Data transformation implementation. The last threat is related to the data transformation implementation. Although the mutation operators are well-designed, there still exists a margin of error in identifying entities and relationships within the input text. To overcome this issue, we use both the Stanza and the Bert models to enhance the precision of entity and relationship identification.
@@ 6. Related Work
This section covers the synthesis method employed by T2I software for text-to-image conversion and discusses the evaluation metrics commonly used in this domain. Following this, we introduce various AI testing methods that are relevant to our study.
6.1. Text-to-Image Synthesis and Evaluation
With the advancement of deep learning, particularly the widespread application of large models in recent times, the performance of T2I software has seen a significant leap forward. Over an extended period, the majority of methods are based on Generative Adversarial Network (GAN), including stacked architecture-based (Zhang et al., 2017, 2018; Bodla et al., 2018), attention-based (Xu et al., 2018; Huang et al., 2019; Tan et al., 2019), siamese architecture-based (Yin et al., 2019; Hadsell et al., 2006; Dumoulin et al., 2016) and cycle consistency-based (Nguyen et al., 2017; Qiao et al., 2019; Lao et al., 2019). As diffusion models make significant breakthroughs in generative tasks, researchers increasingly focus their attention on diffusion structure to construct T2I engine (Gu et al., 2022; Saharia et al., 2022; Rombach et al., 2022; Kim et al., 2022). In addition, autoregressive methods utilize time-series models on large datasets, relying on historical data dependencies across various periods show good performance. A prime example is OpenAI’s DALL-E (Ramesh et al., 2021)and Google’s Parti model (Yu et al., 2022a). However, although more complex model structures enhance the generative effects, T2I software still has inevitable issues.
Therefore, researchers have proposed a series of evaluation metrics to reasonably assess the performance of T2I software, which are usually based on image quality and text-image alignment. Inception score (IS) (Salimans et al., 2016) and Fréchet Inception Distance (FIS) (Bynagari, 2019) are two common indicators to leverage the image quality and image diversity based on pre-trained Inception-v3 network (Szegedy et al., 2016). R-precision (RP) (Xu et al., 2018) is usually used to assess the consistency between text and images. In addition to these general metrics, they also have some new metrics. Tan et al. (Dinh et al., 2022) propose several improved metrics based on IS, FID, SOA, and RP. Cho et al. (Cho et al., 2023) recently introduce two novel interpretable/explainable visual programming frameworks for T2I generation.
However, aside from the issue of overfitting inherent in evaluation metrics themselves, T2I software lacks systematic testing methods to assess its robustness. Unlike the previous methods which focus on evaluating the generation quality and text-image alignment, ACTesting aims to utilize the entity-relationship triple to construct the cross-model testing framework for the generation robustness.
@@ 6.2. AI Testing
Due to the scarcity of testing methods for the T2I task, we introduce some relevant testing methods. After achieving success in traditional testing tasks with mutation testing, researchers propose testing methods to test the deep learning systems. Lei et al. (Ma et al., 2018) apply the mutation testing method and specialize it for deep learning systems to assess the quality of test data. David et al. (Berend et al., 2020) conduct a rich empirical study identifying the impact of mutation operators and coverage criteria on the distribution of the generated deep learning test cases. In addition, Simos et al. (Gerasimou et al., 2020) propose a systematic testing methodology, Importance-Driven(IDC), to assess the semantic diversity of a test set. Currently, one of the most prominent areas of interest in AI testing is the testing methods aimed at autonomous driving (Tian et al., 2018; Feng et al., 2020; Li et al., 2022; Gao et al., 2023). AI testing methods also cover other new areas, including speech recognition systems (Ji et al., 2022), question-answering systems (Liu et al., 2022), and image captioning systems (Yu et al., 2022b).
However, up to this point, no testing methods have been specifically proposed for T2I software. Therefore, building upon cross-modal consistency and mutation testing, we propose ACTesting, a black-box testing method specifically designed for T2I software. This approach is applicable to all T2I software and is effective in detecting defects and issues within them.
@@ 7. Conclusion
In this paper, we introduce ACTesting, an automated, black-box approach grounded in metamorphic testing theory to address the challenges in cross-modal testing for Text-to-Image (T2I) software. ACTesting emphasizes maintaining semantic consistency across different modalities and utilizes the entity-relationship triple to encapsulate key semantic information. We develop three mutation operators based on this triple to identify defects in T2I software, focusing on testing the generation robustness of T2I software. Our evaluation of ACTesting involves four T2I software, generating 113,736 synthetic images, using the MS-COCO validation dataset. We employed metrics I-FID, I-IS, and RP for assessing image realism and text-image relevance. Additionally, we introduced error-detection ($Error_{e}$ and $Error_{r}$) and miss-detection ($Miss_{e}$ and $Miss_{r}$) rates. The findings revealed that ACTesting could degrade image quality by 2.9% to 15% and diminish text-image match consistency by 7.5% to 21.1%. Furthermore, our ablation study showcased the individual effectiveness of each operator, highlighting their capability for flexible combination and superior performance in increasing miss-detection rates compared to basic operators.

### v3

@@ 7. Threats to Validity
Test Subject. The selection of T2I software is one of the vital threats to validity. With the rapid development of T2I software, a diverse array of such software is continually emerging, each with varying engines and mechanisms. To alleviate this threat, we employ five commonly used T2I software powered by different core engines.
@@
Test Data Selection. The selection of the source dataset is another threat. Because not all training and validation datasets used by various software are publicly available, there is a potential data domain unfairness. In addition, prompt construction based on the input text is crucial. To alleviate this threat, we use the most commonly used and classic open-source dataset, MS-COCO, as our seed collection to ensure a level of fairness in the experiments. We also do not include any additional description style or scene details in the system prompt to minimize the impact on the testing process.
Evaluation Model Selection. The last threat is related to the deep learning models used during the evaluation process. To overcome potential biases, we use the regularized Inception-v3 model following (Dinh et al., 2022) to avoid the overfitting problem and apply the open-source pre-trained weights of RvlBert and CLIP.
@@ 8. Related Work
This section covers the evaluation method of T2I software and discusses various AI testing methods relevant to our study.
8.1. Text-to-Image Evaluation
Researchers propose a series of evaluation metrics and methods to reasonably assess the performance of T2I software, which are usually based on image quality and text-image alignment. Inception score (IS) (Salimans et al., 2016) and Fréchet Inception Distance (FIS) (Bynagari, 2019) are two common indicators to leverage the image quality and image diversity based on pre-trained Inception-v3 network (Szegedy et al., 2016). R-precision (RP) (Xu et al., 2018) is usually used to assess the consistency between text and images. Dinh et al. (Dinh et al., 2022) propose several improved metrics based on these metrics. Cho et al. (Cho et al., 2023b) introduce two novel interpretable visual programming frameworks for T2I generation. DALL-Eval  (Cho et al., 2023a) measures three visual reasoning skills, including object recognition and counting, and spatial relation understanding to evaluate T2I models. Saxon et al. (Saxon et al., 2024) introduce T2IScoreScore, a dataset for rigorously evaluating t2i prompt faithfulness metrics. Wan et al.  (Wan et al., 2024) propose DoFaiR, a benchmark to measure the balance between diversity interventions and demographic accuracy in Text-to-Image models. FlashEval (Zhao et al., 2024) is proposed recently to increase the evaluation efficiency by selecting the representative subset of the text-image dataset.
@@ 8.2. AI Testing
Due to the scarcity of testing methods for the T2I task, we introduce some relevant testing methods. After achieving success in traditional testing tasks with metamorphic testing, researchers propose testing methods to test the deep learning systems. Xie et al.  (Xie et al., 2020) develop a metamorphic testing approach to assessing and validating unsupervised machine learning systems. Ma et al. (Ma et al., 2018) apply the metamorphic testing method and specialize it for deep learning systems to assess the quality of test data. Berend et al. (Berend et al., 2020) conduct a rich empirical study identifying the impact of mutation operators and coverage criteria on the distribution of the generated deep learning test cases. Chang et al.  (Chang et al., 2024) introduce VEglue, an object-aligned joint erasing approach for testing Visual Entailment (VE) systems. AI testing methods cover various areas, including autonomous driving (Tian et al., 2018; Feng et al., 2020; Li et al., 2022; Gao et al., 2023), speech recognition systems (Ji et al., 2022), question-answering systems (Liu et al., 2022; Xie et al., 2023), machine translation (Zhang et al., 2024; He et al., 2021; Sun et al., 2020), and image captioning systems (Yu et al., 2022b; Xie et al., 2024).
However, aside from the issue of overfitting inherent in evaluation metrics themselves, T2I software lacks systematic testing methods to assess its performance. Unlike the previous methods, ACTesting aims to apply the metamorphic testing theory and adaptability density constraint, utilizing the ER triple to present the cross-modal semantic information. It constructs the mutation input text and treats the MR as detection criteria to reveal hidden errors of tested T2I software.
@@ 9. Conclusion
In this paper, we introduce ACTesting, an automated, black-box approach grounded in metamorphic testing theory to address the challenges in cross-modal testing for Text-to-Image (T2I) software. ACTesting emphasizes maintaining semantic consistency across different modalities and utilizes the ER triple to encapsulate key semantic information. We develop three mutation operators based on metamorphic testing theory and adaptability density constraint to increase test scenarios. Our experiments involve five T2I software, generating 142,170 synthetic images. The findings reveal that ACTesting could degrade image quality by 2.9% to 15% and diminish text-image match consistency by 7.5% to 21.1%. ACTesting identifies the deficiency of the current T2I software in dealing with unrealistic and complex scenarios. In the future, the quality of the T2I software can be improved by retraining it using input text based on our mutation operators. We also plan to extend ACTesting to a broader range of creative generative artificial intelligence software.
10. Acknowledgement
This work is supported partially by the National Natural Science Foundation of China (61932012, 62372228).